Find the equation of the line that passes through the points (-6,2) and (-5,3).
Write your answer in standrard form.

Answer by fractalier(6550)

First find the slope:
m = (3 - 2) / (-5 - (-6))
m = 1 / 1 = 1
Then use the slope intercept form
y - y1 = m(x - x1)
y - 2 = 1(x + 6)
y = x + 8
which in standard form is
-x + y = 8